media%2F240%2F240b35ef-b6ea-4df3-9c7e-74
Answer #1

Given that

initial velocity u=7 m/s

finial velocity v=9 m/s

time t=8 sec

initial position xo=-27 m

now we find the acceleration

acceleration a=9-7/8=2/8=0.25 m/s^2

now we find the position at t=8 sec

X+27=7*8+1/2*0.25*8^2

x=-27+64=37 m